An investor invested a total of $1800 in two mutual funds. One fund earned a 6% profit value while the other earned 4% profit. If the investor's profit was $80, how much was in each mutual fund? The amount invested in the fund that earned 6% was __________. The amount invested in the fund that earned 4% was ________>
Answer by dragonwalker(73)

let us say that the amount invested in the 6% fund is x. Therefore the amount invested in the 4% fund is 1800-x
Let us change the % to a value, so 4% can be written 4/100 = 0.04 and 6% can be written 0.06
Now we multiply the interest rate by the amount invested for each fund which added together equal $80:
0.06(x) + 0.04(1800-x) = 80
solve for x:
0.06x + 72 - 0.04x = 80
0.02x = 8
To find x divide both sides by 0.02:
0.02x/0.02 = 8/0.02
x = 400
So the amount invested at 6% rate was $400 and therefore the amount at 4% was $1400
To check we multiply the amount invested by the interest rate for each and added together we should get $80
0.06(400) + 0.04(1400) = 80 !!!